MsBrains (ms-brains):

PLEASE HELP WITH 5 QUESTIONS! WILL FAN & MEDAL!!

MsBrains (ms-brains):

I THINK IT'S B...

MsBrains (ms-brains):

@Gabylovesyou @DN_20 Please help me!?

OpenStudy (gabylovesyou):

The answer is D... Solve for r by simplifying both sides of the equation, then isolating the variable.

MsBrains (ms-brains):

oh! okay, I got a few more if youre free?

OpenStudy (gabylovesyou):

im in a middle of a test but.. ill go back and forth so yes :)

MsBrains (ms-brains):

I think it's C

MsBrains (ms-brains):

Alright! I attached a problem

OpenStudy (gabylovesyou):

nevermind . hold up I'm thinking... cx

MsBrains (ms-brains):

Alright. xD I'm going to go listen to my mum, brb

OpenStudy (mrcoolguy):

b

OpenStudy (gabylovesyou):

why ? @MrCoolGuy i keep getting c = b-d/a closest one is D. but I'm really not sure on this question hun... @Ms-Brains

OpenStudy (gabylovesyou):

ik u have to Isolate the variable by dividing each side by factors that don't contain the variable.

MsBrains (ms-brains):

@MrCoolGuy is a troll, and Okay @Gabylovesyou I'll try D

OpenStudy (gabylovesyou):

oh god ._.

OpenStudy (gabylovesyou):

im prob wrong ;(

MsBrains (ms-brains):

I think this is D

OpenStudy (gabylovesyou):

D = 3a + 3b (subtract 3a from both sides) D - 3a = 3a - 3a + 3b D - 3a = 3b (divide both sides by 3) (D - 3a)/3 = b or it could also be written like : D/3 - a = b

OpenStudy (gabylovesyou):

sooooooooo A .

MsBrains (ms-brains):

Just got two more!

OpenStudy (gabylovesyou):

A :)

OpenStudy (gabylovesyou):

again..Solve for w by simplifying both sides of the equation, then isolating the variable.

MsBrains (ms-brains):

LAST ONE!!

MsBrains (ms-brains):

@Gabylovesyou I think it's C...

MsBrains (ms-brains):

@Gabylovesyou where'd you goooooo????

MsBrains (ms-brains):

Awh man. on the last one you leave meeee

OpenStudy (anonymous):

ello :)

OpenStudy (gabylovesyou):

@Ms-Brains IM SORRY. lame connection started to suck ... answer is D because 18 * 11 = 198... 198 * 12 = 2,376

MsBrains (ms-brains):

it's okay! xD I guessed it and got a B. @Gabylovesyou